OpenStudy (anonymous):

I need help with part b, I dont understandthe wording. Now that you have helped Bernard design the pool, he presents you with a situation. There are x number of communities that will be designing homes with the pools. The number of homes within each community is six more than the number of communities. The number of hours that will be required to build each home is double the number of homes in each community. a.Since x represents the number of communities, determine the two other expressions used to represent the number of homes for each community and the number of hours required to build

OpenStudy (anonymous):

Let the no. of communities be x No. of homes are six more than no. of communities Therefore let the number of homes be n \[n = 6 +x\] The number of hours that will be required to build each home is double the number of homes in each community. We know that no. of homes in each communities is n. Let the hours required be h Therefore: \[h = 2n = 2(6 + x) = 12 + 2x\] \[=> h= 12 + 2x\] @agb92101 : your question's answered

OpenStudy (anonymous):

No. Of homes n is: \[n = 6 +x\] No.Of hours h required: \[h = 12x +2\]